### v4

# Phonon-driven multipolar dynamics in a spin-orbit coupled Mott insulator

Abstract
Motivated by advances in pump-probe experiments and light-driven phenomena, we theoretically study the impact of pumped and driven phonons in Mott insulators which host multipole moments, thus going beyond conventional dipolar magnetism. As a case study, we examine pseudospin-1/2 Mott insulators hosting quadrupolar and octupolar moments, and investigate the effect of resonantly exciting ${\cal E}_{g}$ phonon modes which couple linearly to the quadrupoles. We show that this leads to multipolar precession, with the backaction resulting in pseudochiral phonon dynamics in the octupolar ordered phase. We further study the impact of a short-time two-phonon drive, showing that it can induce the build-up of octupolar order or even switch its sign on picosecond timescales. Our results are obtained using a Monte Carlo code incorporating phonons, molecular dynamics simulations to numerically integrate the coupled spin-phonon equations of motion, and analytical Floquet theory. Our work shows how driven phonons can probe and control hidden orders in solids.
preprint: APS/123-QED
Introduction.— Shaking electrons or atoms with light can induce remarkable non-equilibrium many body states in solid state crystals as well as ultracold atoms [1, 2, 3, 4, 5, 6, 7, 8]. In recent years, rapid advances in ultrafast and THz optics have led to the ability to resonantly excite specific orbital transitions and phonon modes in solids [9, 10, 11, 12], providing a distinct pathway to control broken symmetry states including ferroelectricity [13], superconductivity [14], and magnetism [15, 10, 16, 17, 18, 19]. In dipolar magnets, Einstein phonons can strongly impact the spin-spin exchange interaction, stabilizing unconventional equilibrium magnetic orders in geometrically frustrated lattices such as kagome, triangular, or pyrochlore systems [20, 21, 22, 23]. In addition to equilibrium orders, recent experimental and theoretical work has shown that coherent phonon driving can produce fascinating non-equilibrium Floquet dynamics in spin-phonon coupled dipolar magnets. These include anomalously large effective magnetic fields from driven chiral phonons [24, 25, 26, 27, 28, 29, 30], tuning of magnetic orders and magnon band topology [31, 10, 16, 32], frustration breaking in dimerized magnets [33], nonequilibrium phase transitions [34], and high harmonic generation [35]. In addition, this has motivated the study of many-body spin-phonon dynamics and Floquet-phonon physics in layered van der Waals materials [36, 37, 38, 39].
Going beyond dipolar spin systems, there is a rich set of issues one encounters for multipolar degrees of freedom which are found in solids with strong spin-orbit coupling. Unlike dipole moments, higher order multipoles such as quadrupoles and octupoles are associated with spin-orbital entangled charge or current distributions which do not simply couple to conventional probes; for this reason, multipolar broken symmetries, found in diverse materials, are hard to detect and generically termed ‘hidden orders’ [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Quadrupolar moments, however, being even under time-reversal symmetry, can linearly couple to phonons, leading to significant differences in the equilibrium ordering and excitations of multipoles when compared with magnetic dipoles. This is exemplified by recent discoveries of unusual multipolar symmetry breaking and lattice displacements in Ba2MgReO6 [54, 55, 56, 57, 58] and the detection of vibronic modes via resonant inelastic X-ray scattering in K2IrCl6 [59].
Recent experiments have begun to extend optical and THz techniques to study the non-equilibrium dynamics of multipolar orders. For instance, resonantly excited phonon modes in pump-probe and THz experiments have shown evidence of multipolar modes coupled to lattice vibrations [33]. In $\rm{Ca_{2}RuO_{4}}$, THz experiments have revealed a ‘hidden’ quadrupolar-ordered state through the anomalous broadening of coherently driven phonons [60]. These developments call out for a theory of many-body dynamics in pumped and driven multipolar quantum materials as a route to probe and control hidden orders.
In this Letter, we fill this important gap by developing a simulation tool to study multipolar orders and their semiclassical dynamics including the crucial effect of coupling to Einstein phonons with dissipation. This method combines large scale equilibrium spin-phonon Monte Carlo (MC) simulations with spin-phonon molecular dynamics (MD) obtained by integrating their coupled equations of motion. The key idea underlying our work is that a phonon mode coupled to a quadrupole acts as a transverse field to all other orthogonal multipoles, just as for a dipolar spin, a magnetic field along $S_{z}$ acts as a transverse field to $(S_{x},S_{y})$. For multipolar materials, kicking the phonon thus engenders nontrivial precessional dynamics in multipolar space which we leverage as a route to unearth hidden orders.
While our approach is broadly applicable to many multipolar materials, we illustrate the power of this tool by focusing on a recently discovered paradigmatic example in Mott insulators, the osmate double perovskites [61, 62, 63, 53, 64, 65] which display competing octupolar and quadrupolar ordering tendencies [62, 63, 65]. Our work shows how the presence of phonons, resonantly excited by either pump pulses or Floquet driving, can influence the dynamics of the multipolar moments and enable the detection, build-up, and ultrafast switching of exotic octupolar order.
Model.— The simplest example of multipolar moments arise in materials with non-Kramers doublets, local pseudospin-1/2 degrees of freedom whose doublet degeneracy is protected by crystalline point group symmetries rather than just time-reversal symmetry as in the case of Kramers doublets. Remarkably, non-Kramers doublets are found in a diverse array of quantum materials including: quantum spin-ice compounds such as Pr2Zr2O7, Pr2Sn2O7 [66, 67, 68, 69], Kondo lattice intermetallics PrTi2Al20, PrV2Al20 [70, 71, 49, 72, 50], Mott insulating double perovskites $\rm{Ba_{2}CaOsO_{6}}$, $\rm{Ba_{2}MgOsO_{6}}$, $\rm{Ba_{2}ZnOsO_{6}}$ [61, 62, 53], and vacancy ordered double perovskites Cs2WCl6 and Rb2WCl6 [73, 74].
We focus here on non-Kramers doublets with time-reversal even (${\cal T}$-even) quadrupolar moments represented by pseudospin-1/2 Pauli operators $(\tau_{x},\tau_{z})$ and a ${\cal T}$-odd octupole represented by $\tau_{y}$ as realized in the double perovskites and the Kondo intermetallics. In the double perovskites, these pseudospins live on a face-centered cubic (fcc) lattice (see End Matter), and descend from splitting a $J=2$ angular momentum multiplet, with $\tau_{x}\propto(J_{x}^{2}-J_{y}^{2})$, $\tau_{z}\propto(3J_{z}^{2}-J(J+1))$, and $\tau_{y}\propto\overline{J_{x}J_{y}J_{z}}$, with bar denoting operator symmetrization.
Symmetries dictate the following fcc lattice Hamiltonian describing nearest-neighbor pseudospin interactions
$\!\!\mathfrak{h}_{\rm sp}\!$ $=$ $\!\!\!\sum_{\langle{\bf{r}},{\bf{r}}^{\prime}\rangle}\!\!\!\left[-J_{0}\tau_{{\bf{r}}y}\tau_{{\bf{r}}^{\prime}y}\!+\!J_{1}\left(\cos^{2}\!\phi_{{\bf{r}}{\bf{r}}^{\prime}}\!+\!\gamma\sin^{2}\!\phi_{{\bf{r}}{\bf{r}}^{\prime}}\right)\tau_{{\bf{r}}z}\tau_{{\bf{r}}^{\prime}z}\right.$ (1)
$+$ $J_{1}\left(1-\gamma\right)\sin\phi_{{\bf{r}}{\bf{r}}^{\prime}}\cos\phi_{{\bf{r}}{\bf{r}}^{\prime}}\left(\tau_{{\bf{r}}x}\tau_{{\bf{r}}^{\prime}z}+\tau_{{\bf{r}}z}\tau_{{\bf{r}}^{\prime}x}\right)$
$+$ $\left.J_{1}\left(\sin^{2}\phi_{{\bf{r}}{\bf{r}}^{\prime}}+\gamma\cos^{2}\phi_{{\bf{r}}{\bf{r}}^{\prime}}\right)\tau_{{\bf{r}}x}\tau_{{\bf{r}}^{\prime}x}\right],$
where $\phi_{ij}=\{0,2\pi/3,4\pi/3\}$ correspond to nearest neighbors $\langle{\bf{r}}{\bf{r}}^{\prime}\rangle$ in the $\{XY,YZ,ZX\}$ planes. The phase diagram of this model depends on the dimensionless ratios of exchange couplings via $J_{1}/J_{0}$ and $\gamma$ [75, 63].
Based on density functional theory (DFT) and dynamical mean field theory for Ba2CoOsO6 [76], we set $\gamma\!\approx\!-0.4$ and $J_{1}/J_{0}\!\approx\!0.5$ for the pseudospin Hamiltonian; this places us in the ferro-octupolar phase for $T/J_{0}\lesssim 3.2$ as determined from classical Monte Carlo simulations [63]. A microscopic calculation of the octupolar exchange yields $J_{0}\approx 1$ meV, which leads to $T_{\rm oct}\approx 40$ K, in good agreement with experiments on Ba2CoOsO6 [61, 63, 53]. To construct a general phase diagram, we fix $\gamma$ and vary $J_{1}/J_{0}$ and $T/J_{0}$ to access both ferro-octupolar and quadrupolar ordered phases.
Next, we consider cubic symmetry ${\cal E}_{g}$ Einstein phonon modes at each lattice site with $x\equiv x^{2}\!-\!y^{2}$, $z\equiv 3z^{2}\!-\!r^{2}$, with the phonon Hamiltonian
$\mathfrak{h}_{\rm{ph}}=\sum_{{\bf{r}},\alpha=x,z}\left[\frac{1}{2m}\mathfrak{p}_{\alpha}^{2}({\bf{r}})+\frac{1}{2}m\omega^{2}\mathfrak{q}_{\alpha}^{2}({\bf{r}})\right]$ (2)
where $(\mathfrak{q}_{\alpha},\mathfrak{p}_{\alpha})$ are the normal mode coordinates and momenta, and $\omega,m$ represent the frequency and mass, respectively. On symmetry grounds, the ${\cal T}$-odd octupole will not linearly couple to phonons, however the $(\tau_{x},\tau_{z})$ quadrupoles can couple linearly to the phonon modes. The symmetry allowed spin-phonon coupling is given by
$\mathfrak{h}_{\rm sp-ph}=-g\sum_{\bf{r}}\left[\mathfrak{q}_{x}({\bf{r}})\tau_{{\bf{r}}x}+\mathfrak{q}_{z}({\bf{r}})\tau_{{\bf{r}}z}\right]$ (3)
where $g$ is the strength of the spin-phonon coupling.
Figure 1: (a) Monte Carlo phase diagram of the Hamiltonian in Eqs. 1, 2 and 3 as a function of $T/J_{0}$ and $J_{1}/J_{0}$ (for fixed $\gamma=-0.4$). Colors in the ordered phases depict strength of the ferro-octupolar order $\langle\tau_{y}\rangle$ (green) and antiferro-quadrupolar (AFQ) order (purple); insets depict the AFQ and ferro-octupolar orders on two adjacent layers of the fcc lattice (top layer: bold/large, bottom layer: light/small) with orbital shapes and colors respectively showing the charge density and the sign, and the ‘+’ for AFQ indicating a superposition of the two orbital orders. (b) and (c) Phonon dynamics in pump-probe simulations induced by initial momentum kick for $Q_{x}$ phonon for $J_{1}/J_{0}=1.3$ and phonon damping constant $0.072\pi\eta/\omega=$ 0.07 [77, 78], averaged over 160 different initial conditions. (b) $T=0.1J_{0}<T_{\rm oct}$, single domain ferro-octupolar; (c) $T=5.2J_{0}>T_{\rm{oct}}$, paramagnet. For $T<T_{\rm{oct}}$, we observe coherent energy transfer between $Q_{x}$ and $Q_{z}$ phonons, which is absent for $T>T_{\rm{oct}}$. (d) Comparison showing qualitative agreement between octupolar order parameter $M_{y}=\langle\tau_{y}\rangle$ (blue dots) and $Q_{z}$-$Q_{x}$ cross-amplitude, i.e. maximum $Q_{z}$ response to $Q_{x}$ pump (orange dots) as shown in panels (b) and (c); both curves are scaled by their lowest $T$ data point. (e) Schematic of energy exchange between the $(Q_{x},Q_{z})$ phonons in the ferro-octupolar phase through pseudospin precession and quadrupole-phonon coupling.
Equilibrium phase diagram.— We begin by constructing the equilibrium phase diagram of the full model $H=H_{\rm sp}+H_{\rm ph}+H_{\rm sp-ph}$ in the presence of coupling to the cubic ${\cal E}_{g}$ phonon modes. For this purpose, it is useful to define a length scale $a\!=\!0.01a_{0}$, where $a_{0}\!\approx\!4$ is the cubic lattice constant [79], and a normalized phonon amplitude $Q_{\alpha}=\mathfrak{q}_{\alpha}/a$; in these units, $Q_{\alpha}=1$ implies an experimentally accessible $1\%$ atomic displacement, relative to the lattice constant [15]. The normalized conjugate momenta are $P_{\alpha}=a\mathfrak{p}_{\alpha}$. Measuring energy in units of $J_{0}$ yields dimensionless Hamiltonians,
$H_{\rm sp}$ $=$ $\mathfrak{h}_{\rm sp}/J_{0}$ (4)
$H_{\rm ph}$ $=$ $\!\!\!\sum_{{\bf{r}},\alpha=x,z}\left[\frac{P_{\alpha}^{2}({\bf{r}})}{2M}+\frac{1}{2}M\Omega^{2}Q_{\alpha}^{2}({\bf{r}})\right],$ (5)
$H_{\rm sp-ph}$ $=$ $-\lambda\sum_{\bf{r}}\left[Q_{x}({\bf{r}})\tau_{{\bf{r}}x}+Q_{z}({\bf{r}})\tau_{{\bf{r}}z}\right]$ (6)
where the dimensionless coefficients in Eqn. 5, 6 are $1/2M=\hbar^{2}/2mJ_{0}a^{2}$, $\Omega=\hbar\omega/J_{0}$, and $\lambda=ga/J_{0}$. This choice of energy units corresponds to measuring time in units of $\hbar/J_{0}$ ($\approx\!0.6$ picosecond). We set $m=m_{\rm oxygen}\approx 2.6\times 10^{-26}\rm kg$, since we are simulating ${\cal E}_{g}$ phonon modes where only the oxygen atoms have finite displacements, and we set $\hbar\omega\approx 80$ meV as a typical ${\cal E}_{g}$ phonon mode frequency; using this, $1/2M\approx 80$, and $\Omega\approx 80$. From DFT, a $\sim\!1\%$ atomic displacement leads to a $\sim\!5$ meV splitting of the non-Kramers doublet [53]; this implies $g=5{\rm meV}/a$, so the dimensionless spin-phonon coupling $\lambda\approx 5$.
Fig. 1(a) displays the phase diagram obtained from our MC simulations (on system size with $10^{3}$ sites) as we vary the temperature $T/J_{0}$ and quadrupolar exchange $J_{1}/J_{0}$, effectively sampling spin and phonon configurations from thermal populations. We find a high temperature paramagnetic phase, a low temperature ferro-octupolar phase for $J_{1}/J_{0}\!\lesssim\!1.8$ and an antiferro-quadrupolar (AFQ) phase with wavevector $(\pi,\pi,0)$ (and equivalent) for $J_{1}/J_{0}\!\gtrsim\!1.8$. At low $T$, phonon displacements are absent in the ferro-octupolar phase, while the AFQ phase shows small displacements $Q_{\alpha}\!\sim\!0.1$ which track the local quadrupole order. While $Q_{\alpha}({\bf{r}})\neq 0$ at each site, the sign of $Q_{\alpha}({\bf{r}})$ alternates from layer to layer (see Fig. 1). Comparable staggered lattice displacements, $\sim\!0.1\%$, have been found in the AFQ phase of the $j=3/2$ Mott insulator Ba2MgReO6 [56, 58, 54]. We next explore the out-of-equilibrium pseudospin-phonon dynamics.
Detection of octupolar order. — In pump-probe experiments [60], an incoming pump beam can excite the even-parity cubic ${\cal E}_{g}$ phonon modes via a two-photon Raman transition or via non-linear phononics, with the subsequent dynamics providing useful information about broken symmetries. To simulate pump-probe dynamics, we use equilibriated MC spin-phonon configurations as initial conditions, impulsively excite the $Q_{x}$ phonon, and time-evolve the coupled spin and phonon degrees of freedom; see End Matter for simulation details.
Fig. 1(b) shows the time evolution of the phonon displacements $Q_{x}(t)$ and $Q_{z}(t)$ when we give an impulsive kick to the $Q_{x}$ phonon at time $t=0$ in the ferro-octupolar and paramagnetic phases. The rapid oscillations of the $Q_{x},Q_{z}$ modes at the phonon frequency are not visible in the plot, but are shown in the Supplementary Material (SM) [80]. Remarkably, we see that octupolar broken symmetry for $T<T_{\rm oct}$ mediates coherent energy transfer from the $Q_{x}$ to $Q_{z}$ mode on a time scale $t\!\approx\!5\hbar/J_{0}$ ($\approx 3$ ps). Such energy transfer is absent due to rapid dephasing in the paramagnet ($T\!>\!T_{\rm oct}$) or in the quadrupolar phase; this is shown in Fig. 1(c) for the paramagnet.
As shown in Fig. 1(d), the maximum cross-amplitude, i.e. the highest amplitude of $Q_{z}$ oscillations for an impulsive kick to the $Q_{x}$ phonon, serves as a direct probe of the octupolar order parameter. We understand this as follows (see also SM [80]). The octupolar order leads to an average internal field along $\tau_{y}$, while the pumped phonon tips the initial pseudospin state away from $\tau_{y}$. This leads to coherent pseudospin precession around the $\tau_{y}$-axis, serving as a conduit to transfer energy from $Q_{x}$ to $Q_{z}$ and back; see Fig. 1(e). This protocol could be realized using a stimulated Raman scattering mechanism [81] to pump one phonon mode followed by time-resolved X-ray diffraction (trXRD) to probe time-dependent atomic positions for extracting the amplitudes of both phonon modes as a function of time delay. Ref. [82] identifies deformations of order $\sim 3\times 10^{-2}$Å for a bond of length $\sim 2.67$Å, making feasible the detection of the $\sim 1\%$ distortions discussed in our work. The maximum amplitude of the $Q_{z}$ phonon mode, a signature of octupolar order, is revealed on timescales $t\sim 2\hbar/J_{0}$ (about $1.3\rm ps$), which is reasonable for trXRD experiments.
Dynamical build-up of octupolar order. — We next explore how a simultaneous drive of both ${\cal E}_{g}$ phonon modes can be used to control octupolar ordering. We consider $H_{\rm drive}^{\rm ph}\!=\!-A(t)\sum_{\bf{r}}[Q_{x}({\bf{r}})\cos(\Omega t)\!+\!Q_{z}({\bf{r}})\cos(\Omega t-\phi)]$ which describes a coherent drive of Raman-active phonons as a result of terahertz sum-frequency excitation; the drive strength $A(t)$ is thus proportional to the square of the electric field [77, 78]. We implement a short-pulse drive amplitude $A(t)\!=\!A$ for $t\!<\!3\hbar/J_{0}$ and $A(t)\!=\!0$ otherwise. Fig. 2(a) shows the effect of this drive in the paramagnetic phase, averaged over 320 initial MC configurations, for phases $\phi=(0,\pm\pi/4,\pm\pi/2)$. For a drive strength which leads to a peak phonon amplitude $Q_{0}\sim 0.75$, we find that, depending on the relative phase, the drive pulse can induce transient octupolar order $\langle\tau_{y}\rangle$ of either sign, which vanishes at long times. Previous work has shown, using effective rate equations for single-site models, that chiral driving of phonons can induce effective fields for dipolar $4f$ magnets [27, 83, 28]. Our work provides a nontrivial generalization of this idea to multipolar orders on the lattice. To understand this physics, we note that the phonons couple linearly to the quadrupoles, motivating a proxy Hamiltonian $H^{\rm sp}_{\rm drive}=-A_{\rm eff}\sum_{\bf{r}}(\tau_{{\bf{r}}x}\cos(\Omega t)+\tau_{{\bf{r}}z}\cos(\Omega t-\phi))$ which directly drives quadrupoles. The effective field $A_{\rm eff}=\lambda Q_{0}$ with $Q_{0}$ being the peak phonon amplitude (which depends on the drive strength $A$). We study this using a Floquet-Magnus expansion [80] since $\Omega=\hbar\omega/J_{0}\gg 1$, i.e. the phonon dynamics is much faster than pseudospins [5, 3, 84]. In standard notation,
$\!H^{\rm sp}_{\rm drive,0}$ $=$ $0;~{}~{}~{}H^{\rm sp}_{\rm drive,\pm}\!=-\frac{\lambda Q_{0}}{2}\!\sum_{\bf{r}}(\tau_{{\bf{r}}x}\!+\!e^{\mp i\phi}\tau_{{\bf{r}}z})$ (7)
$\!\!\!H^{\rm sp,eff}_{\rm drive}\!$ $=$ $\frac{1}{\Omega}[H^{\rm sp}_{\rm drive,+},H^{\rm sp}_{\rm drive,-}]\!=\frac{\lambda^{2}Q_{0}^{2}}{\Omega}\sin\phi\!\sum_{\bf{r}}\tau_{{\bf{r}}y}$ (8)
This phonon-induced effective field in $H^{\rm sp,eff}_{\rm drive}$, $(-\lambda^{2}Q_{0}^{2}/\Omega)\sin\phi$, explains the induced $\phi$-dependent octupolar order for $T\!>\!T_{\rm oct}$ shown in Fig. 2(a).
Figure 2: Dynamics of $\langle\tau_{y}\rangle$ in the presence of short-time phase coherent two-mode phonon drive with weak dissipation. The drive $H_{\rm drive}^{\rm ph}$ is turned on from $t=0\hbar/J_{0}$ to $t=3\hbar/J_{0}$ and then turned off. (a) In the paramagnet $T\!=\!1.1T_{\rm oct}$, with drive strength $A\!=\!0.5$ (peak phonon amplitude $Q_{0}\!\approx\!0.75$) and various phase differences $\phi\in[-\pi/2,\pi/2]$, showing that driven phonons can select the sign of generated octupolar order. (b) In the ferro-octupolar ordered phase $T\!=\!0.8T_{\rm oct}$, with varying drive strength $A$, which induces a peak phonon amplitude $Q_{0}$ between $2-3$. With a suitable phase choice $\phi\!=\!\pi/2$, it can be used to decrease the octupolar order parameter $\langle\tau_{y}\rangle$, and even switch from positive to negative. Inset: Phonon coordinate $Q_{x}(t)$ for $A\!=\!2$, showing its amplitude build-up and eventual dissipation after the drive is switched off.
Dynamical switching of octupolar order.— We next ask if this pulsed phonon drive can also be used to switch the sign of octupolar order for $T\!<\!T_{\rm oct}$. Fig. 2(b) shows the impact of the two-phonon coherent drive for $\phi=\pi/2$ at $T/T_{\rm oct}=0.8$. Using a drive which leads to a peak phonon amplitude $Q_{0}\!\sim\!1$-$3$ (i.e., $\sim\!1$-$3\%$ atomic displacement), we find that octupolar order in the positive domain is suppressed or can even switch sign, depending on the drive amplitude while the octupolar order in negative domains (not shown) is nearly unaffected. Furthermore, the slow evolution of $\langle\tau_{y}\rangle$ at late times indicate that heating effect of pseudospins is weak at these timescales, although additional cooling might be needed in experiments to extract the energy dissipated by the phonon modes. Nonlinear phononics experiments on CoF2 have realized large atomic displacements $\sim 2\%$ of Raman-active modes by driving an infrared-active phonon with $500$ fs pulse of 10MV/cm electric field [15]; such methods could potentially be extended to the osmates explored in this work.
We find that octupolar switching is mediated by flipping of pseudospins at domain wall boundaries by the phonon-induced field $(-\lambda^{2}Q_{0}^{2}/\Omega)\sin\phi$; for $\phi\!=\!\pi/2$, this promotes growth of $\langle\tau_{y}\rangle\!<\!0$ domains. In Fig.3, we depict this time evolution of pseudospins under a two-phonon drive starting with an initial 3D domain wall configuration with periodic boundary conditions, and a thermal population of phonon momenta. We show a single 2D slice through the lattice, with colours indicating positive $\langle\tau_{y}\rangle$ (red) and negative $\langle\tau_{y}\rangle$ (blue) and the black sites denoting the non-magnetic sublattice. One can clearly see the two-phonon drive inducing a shrinking of the $+\langle\tau_{y}\rangle$ domain as a function of time. The bottom right panel of Fig.3 shows $\langle\tau_{y}\rangle$ averaged over the full lattice, evincing growth of octupolar magnetization; we attribute the superposed rapid oscillations to the high-frequency phonons. Additional data and an effective model of this mechanism is given in the SM [80].
Figure 3: Visualization of octupolar order switching for an initial two-domain configuration (assuming periodic boundary conditions). We drive the $Q_{x}$ and $Q_{z}$ phonons with a relative phase $\phi\!=\!\pi/2$ for a short time $t=3\hbar/J_{0}$ ($\sim 1$ ps) and show configurations at representative times. Red and blue squares respectively depict pseudospins with $\langle\tau_{y}\rangle>0$ and $\langle\tau_{y}\rangle<0$. We clearly see the shrinking of the positive (red) domain and growth of the negative (blue) domain. This domain evolution is also reflected in the net octupolar magnetization $\langle\tau_{y}\rangle$ averaged over the whole lattice (bottom right panel) which is initially zero but evolves to negative values.
Recent remarkable experiments [85] have shown that one can unambiguously diagnose the presence and, significantly, sign of ferro-octupolar order using rotational anisotropy second harmonic generation (SHG). The octupolar order is characterized through interference between the time reversal even and odd channels, with the sign of the octupolar order determining the positions of the measured intensity maxima [85]. We propose a tr-SHG experiment which compares the position of intensity maxima of equilibrium versus the driven system at time $t\sim 3.0\hbar/J_{0}$ ($\sim 2.0\rm ps$) to diagnose the theoretically predicted build-up and switching of octupolar order. Additional plots showing dependence of our results on the phonon damping parameter $\eta$ are given in the SM [80].
Summary.— We have developed computational tools to simulate the equilibrium phases and nonequilibrium dynamics of multipolar degrees of freedom coupled to dissipative Einstein phonons in multipolar Mott insulators. Our equilibrium phase diagram for the model hosts ferro-octupolar and quadrupolar orders at low temperature, as partly described by previous work, but additionally shows that the quadrupolar order is accompanied by weak lattice displacements. We have also shown how pump-probe techniques and two-mode phonon drive can uncover unique dynamical signatures and enable the ultrafast control of the ferro-octupolar state. This study opens up distinct directions in the detection and control of complex hidden orders in quantum materials.
This research was funded by an NSERC Discovery Grant (AP), an Ontario Graduate Scholarship (KH), and an NSERC CGS-D fellowship (RS). We thank Sreekar Voleti, David Hsieh, Prashant Padmanabhan, Emily Zhang, William Bateman-Hemphill, and Cristian Batista for extremely helpful discussions. Numerical computations were performed on the Niagara supercomputer at the SciNet HPC Consortium and the Digital Research Alliance of Canada. Data for the figures in this work is available, see [86].